# Computational lexical analysis of Flamenco genres

### v1

@@ C.4 Malagueñas
In Fig. 21 we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 22 we show the complete $\log P\left(w|C\right)$ distribution for malagueñas.
Figure 21: The 15 words with highest $\log P\left(w|C\right)$ for malagueñas.
Figure 22: Complete $\log P\left(w|C\right)$ distribution for malagueñas over 100 trainings. We mark the threshold used to calculate the number of essential words with a red dot (refer to Appendix B.2).
The essential words for malagueñas in descending order of $P\left(w|C\right)$ are the following:
muerte ’death’, más ’more’, quiero ’I want/love’, querer ’to love’, Dios ’God’, flor ’flower’, quererte ’to love you’, madre ’mother’, persona ’person’, amaba ’I loved’, delirio ’delirium’, Corte ’Court’, mío ’mine’, matando ’killing’, vivir ’to live’.
@@ C.5 Seguiriyas
In Fig. 8c) we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 23 we show the complete $\log P\left(w|C\right)$ distribution for seguiriyas.
The essential words for seguiriyas in descending order of $P\left(w|C\right)$ are the following:
@@
madre ’mother’, Dios ’God’, corazón ’heart’, alma ’soul’, más ’more’, mía ’mine (fem. sing.)’, Que/Qué ’Which/What’, mal ’harm’, grandes ’big (pl.)’, mío ’mine’, libertad ’freedom’, andar ’to walk’, encuentro ’I find’, doy ’I give’, desgracia ’misfortune’, Santiago ’Santiago’, hospital ’hospital’, calor ’heat’, veo ’I see’, rincones ’corners’, pido ’I ask’, llorando ’crying’, fatigas ’fatigues’, quiero ’I want/love’, muero ’I die’, muerte ’death’, voy ’I go’, pago ’I pay’, Siempre ’Always’, Por ’For’, paso ’step’, aliviara ’he/she/I would alleviate’, penas ’pains, sorrows’, duquelas ’pains, sorrows (gipsy lang.)’, noche ’night’, padre ’father’, vida ’life’, reniego ’I renounce’, niños ’children’, cuerpo ’body’, compañerita ’little companion (fem.)’, día ’day’, clavito ’little clove’, adelante ’forward’, clavo ’clove’, pena ’pain, sorrow’, compañera ’companion (fem.)’, dolores ’pains’, puerta ’door’, hecho ’done’, Santa Ana ’Saint Anne’, dos ’two’, nadie ’nobody’, rodillas ’knees’, atrás ’behind’, apelación ’appeal’, rogue ’I beg’, viene ’he/she comes’, Como ’Like, I eat’, grande ’big’, muera ’I die’, caben ’they fit’, Si ’If’, llamarme ’to call me’, sabe ’he/she knows’, lengua ’tongue’, huele ’it smells’, dolor ’pain’, voces ’voices’, No ’No’, pierda ’I lose’, solo ’alone’, perdón ’forgiveness’, vera ’side, proximity’, llamo ’I call’, Yo ’I’, muerto ’dead’, canela ’cinnamon’, daba ’he/she gave’, dieron ’they gave’, encargo ’order’, tan ’so’, Él ’He’, doce ’twelve’, saber ’to know’, llevo ’I carry’, supiera ’he/she knew’, hueles ’you smell’, mando ’I send, I command’, cielo ’sky’, pague ’I/he/she pay’, bien ’well’, mala ’bad (fem.)’, limosna ’alms’, malita ’poor thing (fem.)’, tierra ’land’, vienes ’you come’, sueño ’dream’, médico ’doctor’, horita ’little hour’, dormido ’asleep’, camelo ’I want/love (gipsy lang.)’, eternidad ’eternity’, aquella ’that (fem.)’, hora ’hour’, pasos ’steps’, Te ’You’, había ’there was’, dobles ’double (pl.)’, llores ’you cry’, echo ’I throw’, distinguir ’to distinguish’, San ’Saint’, terelo ’I have’, morir ’to die’, van ’they go’, ir ’to go’, siempre ’always’, dile ’tell him/her’, quemando ’burning’, ver ’to see’, dar ’to give’, mujer ’woman’, Me ’Me’, medio ’half’, siete ’seven’, pasitos ’little steps’, doctor ’doctor’, metida ’inserted (fem. sing.)’, llorar ’to cry’, quieres ’you want’, ladrón ’thief’, remedio ’remedy’, manos ’hands’, señaladitos ’pointed’, ahora ’now’, vine ’I came’, mandé ’I sent’, gente ’people’, solito ’alone’, De ’Of’, calle ’street’, Cuco ’Cuco’, sangre ’blood’, duermes ’you sleep’, vas ’you go’, respondía ’he/she answered’, rodar ’to roll’, negrito ’little black one’, pidiendo ’asking’, días ’days’, malito ’little sick one’, acabo ’I finish’, Eran ’they were’, culpa ’fault’, hermano ’brother’, fatiguitas ’little fatigues’, pasando ’passing’, mundo ’world’, conozco ’I know’, conocí ’I knew’, basta ’enough’, hija ’daughter’, sobra ’it is enough’, Contemplarme ’To contemplate me’, ojitos ’little eyes’, llamar ’to call’, trenzas ’braids’, causas ’causes’, así ’like this’, gitana ’gypsy (fem. sing.)’, muriendo ’dying’, carta ’letter’.
@@ C.6 Soleá
In Fig. 24 we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 25 we show the complete $\log P\left(w|C\right)$ distribution for soleá.
Figure 24: The 15 words with highest $\log P\left(w|C\right)$ for soleá.
@@
Figure 25: Complete $\log P\left(w|C\right)$ distribution for soleá over 100 trainings. We mark the threshold used to calculate the number of essential words with a red dot (refer to Appendix B.2).
The essential words for soleá in descending order of $P\left(w|C\right)$ are the following:
@@
Dios ’God’, quiero ’I want/love’, querer ’to love’, nadie ’nobody’, Yo ’I’, madre ’mother’, pena ’pain, sorrow’, voy ’I go’, tan ’so’, No ’No’, corazón ’heart’, Que/Qué ’Which/What’, más ’more’, Por ’For’, cara ’face’, quiera ’I/he/she wants’, mundo ’world’, Undebel ’the devil (gipsy lang.)’, mía ’mine (fem. sing.)’, mal ’bad’, ahora ’now’, ver ’to see’, había ’there was’, gitana ’gypsy (fem.)’, hago ’I do’, dos ’two’, lengua ’tongue’, contigo ’with you’, alma ’soul’, cuerpo ’body’, sol ’sun’, Tu ’Your’, gusto ’taste’, puerta ’door’, puedo ’I can’, quieres ’you want’, vera ’side, proximity’, Si ’If’, tierra ’land’, mala ’bad’, quería ’I/he/she wanted’, Me ’Me’, ser ’to be’, vez ’time’, vas ’you go’, muerte ’death’, mío ’mine’, centro ’center’, nunca ’never’, ojos ’eyes’, metida ’inserted (fem. sing.)’, aire ’air’, penas ’pains, sorrows’, presente ’present’, tiempo ’time’, gente ’people’, prima ’cousin (fem.)’, conmigo ’with me’, calle ’street’, hecho ’done’, hablar ’to talk’, casa ’house’, carnes ’meats’, piedras ’stones’, va ’goes’, vida ’life’, camelo ’I want/love (gipsy lang.)’, día ’day’, creo ’I believe’, sentido ’sense’, bien ’well’, encuentro ’I find, encounter’, compañera ’companion (fem.)’, fatigas ’fatigues’, loco ’crazy’, perdido ’lost’, paso ’step’, tenía ’I/he/she had’, perdone ’I/he/she forgive/s’, sale ’he/she leaves’, verás ’you will see’, flamenca ’flamenca’, viene ’he/she comes’, grande ’big’, quisiera ’I would like/want’, culpa ’fault’, piedra ’stone’, salud ’health’, Él ’He’, echaría ’he/she would throw’, remedio ’remedy’, aunque ’although’, Cuando ’When’, querido ’dear’, loca ’crazy (fem. sing.)’, ciencia ’science’, voces ’voices’, alegría ’joy’, cielo ’sky’, tenerte ’to have you’, ojitos ’little eyes’, persona ’person’, san ’saint’, conoces ’you know’, hablo ’I speak’, colorado ’colored’, pido ’I ask’, campanas ’bells’, morir ’to die’, ’I know’, da ’gives’, malina ’evil (fem. sing.)’, veo ’I see’, camino ’path’, entonces ’then’, Señor ’Lord’, duquelas ’pains, sorrows (gipsy lang.)’, mira ’he/she looks’, pensar ’to think’, perder ’to lose’, pasa ’it happens’, verdad ’truth’, noche ’night’, tener ’to have’, Como ’Like’, Dices ’You say’, quedo ’I stay’, mañana ’morning’, pasar ’to pass’, Tengo ’I have’, quieras ’you want’, vienes ’you come’, vivir ’to live’, Válgame ’Help me’, lleno ’full’, infierno ’hell’, intención ’intention’, muerto ’dead’, Para ’For’, bajabas ’you descended’, llorar ’to cry’, dado ’given’, delante ’in front of’, perdición ’perdition’, borde ’edge’, será ’it will be’, llamo ’I call’, Te ’You’, abrirme ’to open myself’, tres ’three’, padre ’father’, perdón ’forgiveness’, gloria ’glory’, vayas ’you go’, razón ’reason’, pase ’it happens’, murió ’he/she died’, dejado ’left’, eche ’I throw’, perdí ’I lost’, mar ’sea’, vestir ’to dress’, dije ’I said’, La ’The (fem. sing.)’, señalando ’pointing’, queriendo ’wanting/loving’, cuenta ’account’, pasado ’past’, están ’they are’, puesto ’placed’, válgame ’help me’, Con ’With’, ley ’law’, cabecita ’little head’, negra ’black (fem. sing.)’, compañerita ’little companion (fem. sing.)’, llevo ’I carry’, serrana ’mountain woman’, sangre ’blood’, calles ’streets, you shut up’, solamente ’only’, dejarme ’to leave me’, siempre ’always’.
@@ C.7 Tangos
In Fig. 26 we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 27 we show the complete $\log P\left(w|C\right)$ distribution for tangos.
Figure 26: The 15 words with highest $\log P\left(w|C\right)$ for tangos.
Figure 27: Complete $\log P\left(w|C\right)$ distribution for tangos over 100 trainings. We mark the threshold used to calculate the number of essential words with a red dot (refer to Appendix B.2).
The essential words for tangos in descending order of $P\left(w|C\right)$ are the following:
quiero ’I want/love’, más ’more’, madre ’mother’, Que/Qué ’Which/What’, querer ’to love’, quieres ’you want’, Si ’If’, Yo ’I’, pena ’pain’, calle ’street’, voy ’I go’, nadie ’nobody’, cara ’face’, guerra ’war’, gitana ’gypsy (fem.)’, verás ’you will see’, cuerpo ’body’, Tu ’Your’, vienes ’you come’, Santa María ’Saint Mary’, agua ’water’, mira ’you look, he/she looks’, corazón ’heart’, día ’day’, mía ’mine (fem. sing.)’, noche ’night’, bien ’well’, casa ’house’, mías ’mine’, tonta ’silly’, No ’No’, negros ’black (masc. pl.)’, Dios ’God’, quiera ’he/she wants’, serrana ’mountain woman’, gente ’people’, hace ’he/she does’, alegría ’joy’, vez ’time’, pasa ’it happens’, bonita ’beautiful (fem. sing.)’, mañana ’morning’, Triana ’Triana’, Virgen ’Virgin’, contigo ’with you’, Utrera ’Utrera’, Mi ’My’, mundo ’world’, ser ’to be’, puente ’bridge’, encender ’to light’, niño ’child’, Al ’To the’, fatigas ’fatigues’, flores ’flowers’, Humildad ’Humility’, va ’he/she goes’.
@@ C.8 Tientos
In Fig. 28 we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 29 we show the complete $\log P\left(w|C\right)$ distribution for tientos.
Figure 28: The 15 words with highest $\log P\left(w|C\right)$ for tientos.
Figure 29: Complete $\log P\left(w|C\right)$ distribution for tientos over 100 trainings. We mark the threshold used to calculate the number of essential words with a red dot (refer to Appendix B.2).
The essential words for tientos in descending order of $P\left(w|C\right)$ are the following:
querer ’to want/love’, digo ’I say’, más ’more’, Dios ’God’, madre ’mother’, conmigo ’with me’, quiero ’I want/love’, muerte ’death’, fatigas ’fatigues’, hablo ’I speak’, tonta ’silly (fem. sing.)’, carita ’little face’, vida ’life’, vienes ’you come’, tan ’so (comp.)’, dado ’given’, llorar ’to cry’, pocito ’little well’, mía ’mine (fem. sing.)’, grande ’big (sing.)’, hecho ’made (masc. sing.)’, cuerpo ’body’, día ’day’, motivo ’reason’, siento ’I feel’, serrana ’mountain woman’, Yo ’I’, pena ’pain, sorrow’, gitana ’gypsy (fem. sing.)’, veo ’I see’, vera ’side, proximity’, inocente ’innocent (sing.)’, corazón ’heart’, Virgen ’Virgin’, daré ’I will give’, voy ’I go’, bien ’well’, agüita ’little water’, aquel ’that one’, rico ’rich (masc. sing.)’, Si ’If’, verita ’little vera’, Que/Qué ’Which/What’, temblé ’I trembled’, dan ’they give’, verte ’to see you’, escucharlo ’to listen to it’, agua ’water’, parece ’it seems’, mentira ’lie’, vente ’you come’, contigo ’with you’, grandes ’big (pl.)’, dando ’giving’, guerra ’war’.

### v2

@@ C.4 Malagueñas
In Fig. 16(a) we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 16(b) we show the complete $\log P\left(w|C\right)$ distribution for malagueñas. The essential words for malagueñas in descending order of $P\left(w|C\right)$ are the following:
Figure 16: (a) The 15 words with highest $\log P\left(w|C\right)$ for malagueñas, and (b) complete $\log P\left(w|C\right)$ distribution for malagueñas over 100 trainings. We mark the threshold used to calculate the number of essential words with a red dot (refer to Appendix B.2).
muerte ‘death’, más ‘more’, quiero ‘I want/love’, querer ‘to love’, Dios ‘God’, flor ‘flower’, quererte ‘to love you’, madre ‘mother’, persona ‘person’, amaba ‘I loved’, delirio ‘delirium’, Corte ‘Court’, mío ‘mine’, matando ‘killing’, vivir ‘to live’.
@@ C.5 Seguiriyas
In Fig. 6c) we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 17 we show the complete $\log P\left(w|C\right)$ distribution for seguiriyas. The essential words for seguiriyas in descending order of $P\left(w|C\right)$ are the following:
@@
madre ‘mother’, Dios ‘God’, corazón ‘heart’, alma ‘soul’, más ‘more’, mía ‘mine (fem. sing.)’, Que/Qué ‘Which/What’, mal ‘harm’, grandes ‘big (pl.)’, mío ‘mine’, libertad ‘freedom’, andar ‘to walk’, encuentro ‘I find’, doy ‘I give’, desgracia ‘misfortune’, Santiago ‘Santiago’, hospital ‘hospital’, calor ‘heat’, veo ‘I see’, rincones ‘corners’, pido ‘I ask’, llorando ‘crying’, fatigas ‘fatigues’, quiero ‘I want/love’, muero ‘I die’, muerte ‘death’, voy ‘I go’, pago ‘I pay’, Siempre ‘Always’, Por ‘For’, paso ‘step’, aliviara ‘he/she/I would alleviate’, penas ‘pains, sorrows’, duquelas ‘pains, sorrows (gipsy lang.)’, noche ‘night’, padre ‘father’, vida ‘life’, reniego ‘I renounce’, niños ‘children’, cuerpo ‘body’, compañerita ‘little companion (fem.)’, día ‘day’, clavito ‘little clove’, adelante ‘forward’, clavo ‘clove’, pena ‘pain, sorrow’, compañera ‘companion (fem.)’, dolores ‘pains’, puerta ‘door’, hecho ‘done’, Santa Ana ‘Saint Anne’, dos ‘two’, nadie ‘nobody’, rodillas ‘knees’, atrás ‘behind’, apelación ‘appeal’, rogue ‘I beg’, viene ‘he/she comes’, Como ‘Like, I eat’, grande ‘big’, muera ‘I die’, caben ‘they fit’, Si ‘If’, llamarme ‘to call me’, sabe ‘he/she knows’, lengua ‘tongue’, huele ‘it smells’, dolor ‘pain’, voces ‘voices’, No ‘No’, pierda ‘I lose’, solo ‘alone’, perdón ‘forgiveness’, vera ‘side, proximity’, llamo ‘I call’, Yo ‘I’, muerto ‘dead’, canela ‘cinnamon’, daba ‘he/she gave’, dieron ‘they gave’, encargo ‘order’, tan ‘so’, Él ‘He’, doce ‘twelve’, saber ‘to know’, llevo ‘I carry’, supiera ‘he/she knew’, hueles ‘you smell’, mando ‘I send, I command’, cielo ‘sky’, pague ‘I/he/she pay’, bien ‘well’, mala ‘bad (fem.)’, limosna ‘alms’, malita ‘poor thing (fem.)’, tierra ‘land’, vienes ‘you come’, sueño ‘dream’, médico ‘doctor’, horita ‘little hour’, dormido ‘asleep’, camelo ‘I want/love (gipsy lang.)’, eternidad ‘eternity’, aquella ‘that (fem.)’, hora ‘hour’, pasos ‘steps’, Te ‘You’, había ‘there was’, dobles ‘double (pl.)’, llores ‘you cry’, echo ‘I throw’, distinguir ‘to distinguish’, San ‘Saint’, terelo ‘I have’, morir ‘to die’, van ‘they go’, ir ‘to go’, siempre ‘always’, dile ‘tell him/her’, quemando ‘burning’, ver ‘to see’, dar ‘to give’, mujer ‘woman’, Me ‘Me’, medio ‘half’, siete ‘seven’, pasitos ‘little steps’, doctor ‘doctor’, metida ‘inserted (fem. sing.)’, llorar ‘to cry’, quieres ‘you want’, ladrón ‘thief’, remedio ‘remedy’, manos ‘hands’, señaladitos ‘pointed’, ahora ‘now’, vine ‘I came’, mandé ‘I sent’, gente ‘people’, solito ‘alone’, De ‘Of’, calle ‘street’, Cuco ‘Cuco’, sangre ‘blood’, duermes ‘you sleep’, vas ‘you go’, respondía ‘he/she answered’, rodar ‘to roll’, negrito ‘little black one’, pidiendo ‘asking’, días ‘days’, malito ‘little sick one’, acabo ‘I finish’, Eran ‘they were’, culpa ‘fault’, hermano ‘brother’, fatiguitas ‘little fatigues’, pasando ‘passing’, mundo ‘world’, conozco ‘I know’, conocí ‘I knew’, basta ‘enough’, hija ‘daughter’, sobra ‘it is enough’, Contemplarme ‘To contemplate me’, ojitos ‘little eyes’, llamar ‘to call’, trenzas ‘braids’, causas ‘causes’, así ‘like this’, gitana ‘gypsy (fem. sing.)’, muriendo ‘dying’, carta ‘letter’.
@@ C.6 Soleá
In Fig. 18(a) we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 18(b) we show the complete $\log P\left(w|C\right)$ distribution for soleá. The essential words for soleá in descending order of $P\left(w|C\right)$ are the following:
@@
Figure 18: (a) The 15 words with highest $\log P\left(w|C\right)$ for soleá, and (b) complete $\log P\left(w|C\right)$ distribution for soleá over 100 trainings. We mark the threshold used to calculate the number of essential words with a red dot (refer to Appendix B.2).
@@
Dios ‘God’, quiero ‘I want/love’, querer ‘to love’, nadie ‘nobody’, Yo ‘I’, madre ‘mother’, pena ‘pain, sorrow’, voy ‘I go’, tan ‘so’, No ‘No’, corazón ‘heart’, Que/Qué ‘Which/What’, más ‘more’, Por ‘For’, cara ‘face’, quiera ‘I/he/she wants’, mundo ‘world’, Undebel ‘the devil (gipsy lang.)’, mía ‘mine (fem. sing.)’, mal ‘bad’, ahora ‘now’, ver ‘to see’, había ‘there was’, gitana ‘gypsy (fem.)’, hago ‘I do’, dos ‘two’, lengua ‘tongue’, contigo ‘with you’, alma ‘soul’, cuerpo ‘body’, sol ‘sun’, Tu ‘Your’, gusto ‘taste’, puerta ‘door’, puedo ‘I can’, quieres ‘you want’, vera ‘side, proximity’, Si ‘If’, tierra ‘land’, mala ‘bad’, quería ‘I/he/she wanted’, Me ‘Me’, ser ‘to be’, vez ‘time’, vas ‘you go’, muerte ‘death’, mío ‘mine’, centro ‘center’, nunca ‘never’, ojos ‘eyes’, metida ‘inserted (fem. sing.)’, aire ‘air’, penas ‘pains, sorrows’, presente ‘present’, tiempo ‘time’, gente ‘people’, prima ‘cousin (fem.)’, conmigo ‘with me’, calle ‘street’, hecho ‘done’, hablar ‘to talk’, casa ‘house’, carnes ‘meats’, piedras ‘stones’, va ‘goes’, vida ‘life’, camelo ‘I want/love (gipsy lang.)’, día ‘day’, creo ‘I believe’, sentido ‘sense’, bien ‘well’, encuentro ‘I find, encounter’, compañera ‘companion (fem.)’, fatigas ‘fatigues’, loco ‘crazy’, perdido ‘lost’, paso ‘step’, tenía ‘I/he/she had’, perdone ‘I/he/she forgive/s’, sale ‘he/she leaves’, verás ‘you will see’, flamenca ‘flamenca’, viene ‘he/she comes’, grande ‘big’, quisiera ‘I would like/want’, culpa ‘fault’, piedra ‘stone’, salud ‘health’, Él ‘He’, echaría ‘he/she would throw’, remedio ‘remedy’, aunque ‘although’, Cuando ‘When’, querido ‘dear’, loca ‘crazy (fem. sing.)’, ciencia ‘science’, voces ‘voices’, alegría ‘joy’, cielo ‘sky’, tenerte ‘to have you’, ojitos ‘little eyes’, persona ‘person’, san ‘saint’, conoces ‘you know’, hablo ‘I speak’, colorado ‘colored’, pido ‘I ask’, campanas ‘bells’, morir ‘to die’, ‘I know’, da ‘gives’, malina ‘evil (fem. sing.)’, veo ‘I see’, camino ‘path’, entonces ‘then’, Señor ‘Lord’, duquelas ‘pains, sorrows (gipsy lang.)’, mira ‘he/she looks’, pensar ‘to think’, perder ‘to lose’, pasa ‘it happens’, verdad ‘truth’, noche ‘night’, tener ‘to have’, Como ‘Like’, Dices ‘You say’, quedo ‘I stay’, mañana ‘morning’, pasar ‘to pass’, Tengo ‘I have’, quieras ‘you want’, vienes ‘you come’, vivir ‘to live’, Válgame ‘Help me’, lleno ‘full’, infierno ‘hell’, intención ‘intention’, muerto ‘dead’, Para ‘For’, bajabas ‘you descended’, llorar ‘to cry’, dado ‘given’, delante ‘in front of’, perdición ‘perdition’, borde ‘edge’, será ‘it will be’, llamo ‘I call’, Te ‘You’, abrirme ‘to open myself’, tres ‘three’, padre ‘father’, perdón ‘forgiveness’, gloria ‘glory’, vayas ‘you go’, razón ‘reason’, pase ‘it happens’, murió ‘he/she died’, dejado ‘left’, eche ‘I throw’, perdí ‘I lost’, mar ‘sea’, vestir ‘to dress’, dije ‘I said’, La ‘The (fem. sing.)’, señalando ‘pointing’, queriendo ‘wanting/loving’, cuenta ‘account’, pasado ‘past’, están ‘they are’, puesto ‘placed’, válgame ‘help me’, Con ‘With’, ley ‘law’, cabecita ‘little head’, negra ‘black (fem. sing.)’, compañerita ‘little companion (fem. sing.)’, llevo ‘I carry’, serrana ‘mountain woman’, sangre ‘blood’, calles ‘streets, you shut up’, solamente ‘only’, dejarme ‘to leave me’, siempre ‘always’.
@@ C.7 Tangos
In Fig. 19(a) we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 19(b) we show the complete $\log P\left(w|C\right)$ distribution for tangos. The essential words for tangos in descending order of $P\left(w|C\right)$ are the following:
Figure 19: (a) The 15 words with highest $\log P\left(w|C\right)$ for tangos, and (b) complete $\log P\left(w|C\right)$ distribution for tangos over 100 trainings. We mark the threshold used to calculate the number of essential words with a red dot (refer to Appendix B.2).
quiero ‘I want/love’, más ‘more’, madre ‘mother’, Que/Qué ‘Which/What’, querer ‘to love’, quieres ‘you want’, Si ‘If’, Yo ‘I’, pena ‘pain’, calle ‘street’, voy ‘I go’, nadie ‘nobody’, cara ‘face’, guerra ‘war’, gitana ‘gypsy (fem.)’, verás ‘you will see’, cuerpo ‘body’, Tu ‘Your’, vienes ‘you come’, Santa María ‘Saint Mary’, agua ‘water’, mira ‘you look, he/she looks’, corazón ‘heart’, día ‘day’, mía ‘mine (fem. sing.)’, noche ‘night’, bien ‘well’, casa ‘house’, mías ‘mine’, tonta ‘silly’, No ‘No’, negros ‘black (masc. pl.)’, Dios ‘God’, quiera ‘he/she wants’, serrana ‘mountain woman’, gente ‘people’, hace ‘he/she does’, alegría ‘joy’, vez ‘time’, pasa ‘it happens’, bonita ‘beautiful (fem. sing.)’, mañana ‘morning’, Triana ‘Triana’, Virgen ‘Virgin’, contigo ‘with you’, Utrera ‘Utrera’, Mi ‘My’, mundo ‘world’, ser ‘to be’, puente ‘bridge’, encender ‘to light’, niño ‘child’, Al ‘To the’, fatigas ‘fatigues’, flores ‘flowers’, Humildad ‘Humility’, va ‘he/she goes’.
@@ C.8 Tientos
In Fig. 20(a) we show the 15 words with the highest $\log P\left(w|C\right)$, and in Fig. 20(b) we show the complete $\log P\left(w|C\right)$ distribution for tientos. The essential words for tientos in descending order of $P\left(w|C\right)$ are the following:
Figure 20: (a) The 15 words with highest $\log P\left(w|C\right)$ for tientos, and (b) complete $\log P\left(w|C\right)$ distribution for tientos over 100 trainings. We mark the threshold used to calculate the number of essential words with a red dot (refer to Appendix B.2).
querer ‘to want/love’, digo ‘I say’, más ‘more’, Dios ‘God’, madre ‘mother’, conmigo ‘with me’, quiero ‘I want/love’, muerte ‘death’, fatigas ‘fatigues’, hablo ‘I speak’, tonta ‘silly (fem. sing.)’, carita ‘little face’, vida ‘life’, vienes ‘you come’, tan ‘so (comp.)’, dado ‘given’, llorar ‘to cry’, pocito ‘little well’, mía ‘mine (fem. sing.)’, grande ‘big (sing.)’, hecho ‘made (masc. sing.)’, cuerpo ‘body’, día ‘day’, motivo ‘reason’, siento ‘I feel’, serrana ‘mountain woman’, Yo ‘I’, pena ‘pain, sorrow’, gitana ‘gypsy (fem. sing.)’, veo ‘I see’, vera ‘side, proximity’, inocente ‘innocent (sing.)’, corazón ‘heart’, Virgen ‘Virgin’, daré ‘I will give’, voy ‘I go’, bien ‘well’, agüita ‘little water’, aquel ‘that one’, rico ‘rich (masc. sing.)’, Si ‘If’, verita ‘little vera’, Que/Qué ‘Which/What’, temblé ‘I trembled’, dan ‘they give’, verte ‘to see you’, escucharlo ‘to listen to it’, agua ‘water’, parece ‘it seems’, mentira ‘lie’, vente ‘you come’, contigo ‘with you’, grandes ‘big (pl.)’, dando ‘giving’, guerra ‘war’.